OpenStudy (anonymous):

Solve for x: 2 over 5 (x - 4) = 2x.

OpenStudy (anonymous):

What...? Why'd you move the 2 around?

jimthompson5910 (jim_thompson5910):

The equation is \[\Large \frac{2}{5}(x-4) = 2x\] right?

OpenStudy (anonymous):

Yessir

jimthompson5910 (jim_thompson5910):

Multiply both sides by 5 \[\Large \frac{2}{5}(x-4) = 2x\] \[\Large 5*\frac{2}{5}(x-4) = 5*2x\] \[\Large 2(x-4) = 10x\]

jimthompson5910 (jim_thompson5910):

Then distribute \[\Large 2(x-4) = 10x\] \[\Large 2x-8 = 10x\] I'll let you finish up

OpenStudy (anonymous):

I've got no idea... I tried adding 8 to both sides, then dividing, and I tried leaving it and filling in for x.. Is there some mathematical way to find x?

jimthompson5910 (jim_thompson5910):

why not subtract 2x from both sides

jimthompson5910 (jim_thompson5910):

what do you get?

OpenStudy (anonymous):

\[2x-2x - 8 =10x -2x\] \[- 8 =8x \] \[\frac{ - 8 =8x }{ 8 }\] \[x = -1\]

jimthompson5910 (jim_thompson5910):

you got it

jimthompson5910 (jim_thompson5910):

line 3 has odd notation, but that's fine really

OpenStudy (anonymous):

Thank you. I forgot you could subtract the 2x on both sides. I was meant to represent division

OpenStudy (anonymous):

It*

jimthompson5910 (jim_thompson5910):

np